given the following equation U=√X1 X2,. what would be the total effect of a price change when the price of one good reduces by 10% from shs. 30. the income of the consumer is shs. 6000 per month while the price of the good two is 40


1
Expert's answer
2021-12-05T18:56:12-0500

U=√X1 X2

"U=X1^{0.5}X2^{0.5}"


"MU_{X1}=\frac{\delta U}{\delta X1}=0.5X1^{-0.5}X2^{0.5}"


"MU_{X2}=\frac{\delta U}{\delta X2}=0.5X1^{0.5}X2^{-0.5}"


The budget line is;

"P1\times X1 + P2\times X2= I"


"30X1+40X2=6000" ……..equation 1


"X1=\frac{6000-40X2}{30}"


"X2=\frac{6000-30X1}{40}"


When price of good 1 reduces by "10\%" from 30,equation 1 becomes;

"27X1+40X2=6000"


"X1=\frac{6000-40X2}{27}=222.22-1.48X2"


"X2=\frac{6000-27X1}{40}=150-0.675X1"


If the price of good 1 reduces by 10% from Ksh. 30, then the quantity of X1 consumed will increase, and the quantity of X2 consumed will decrease.

But the total effect of a price change will be positive, so the total utility will increase.